Mike was 25 years old when his son was born. Mike's present age is 7 years less than three times his son's present age. How old is Mike now?
Answer by ankor@dixie-net.com(22740)

Let M = Mikes age now
Let S = Son's age now
:
From this statement,"Mike was 25 years old when his son was born." we can say:
S = M - 25
;
Mike's present age is 7 years less than three times his son's present age.
M = 3S - 7
:
How old is Mike now?
Substitute (M-25) for S for S in the above equation and solve for M
M = 3(M-25) - 7
M = 3M - 75 - 7
M - 3M = - 82
-2M = -82
M = -82/-2
M = +41 yrs is Mike's age now
:
:
Find Son's age and check solution in the statement:
S = 41 - 25 = 16
" Mike's present age is 7 years less than three times his son's present age."
41 = 3(16) - 7
41 = 48 - 7; confirms our solution